OpenStudy (anonymous):

f(x)=e^[a(x^2)+bx+1] , find a and b , if f(1)=f(0)=f'(0) .

OpenStudy (jamesj):

Well, f(0) = e^1 and f(1) = e^(a+b+1) hence if f(0) = f(1), it must be that a+b = 0. Now find the other equation for a and b, using f'(x)

OpenStudy (anonymous):

Hmmm.. can you show how?

OpenStudy (anonymous):

You need to find the derivative of the function

OpenStudy (anonymous):

\[f'(x)=(2ax+b)e^{ax^2+bx+1}\]

OpenStudy (jamesj):

Right. Now evaluate that expression at zero, f'(0) and set it equal to the earlier expressions.

OpenStudy (anonymous):

Thank you !